5. What is the apparent weight of a 75 kg astronaut 4200 km from the centre of Earth’s
moon in a space vehicle?
a) Moving at constant velocity
b) Accelerating toward the moon at 2.9 m/s2
1
Expert's answer
2020-05-21T13:04:41-0400

Weight: w = 75kg

Distance: d = 4200km


Part(a)

moving at constant velocity,

"g=\frac{GM_{moon}}{r^2}"


="\frac{6.67×10^{−11}×7.35×10^{22}}{(4.2×10^6)^2}"

="209 m\/s=0.277m\/s^2"


Apparent weight:

=75×0.277=20.775N


(b)

Accelerating towards the moon at

Total force:

=75×2.9=217N


Thus, apparent weight is:

=257−20.775=236.225N


Radially away from the moon.